# Tailoring with Targeted Precision: Edit-Based Agents for Open-Domain Procedure Customization

### v2

@@ Abstract
[…]
@@ 2 Task Setup
In this section, we define a problem formulation and evaluation scheme for procedure customization and use showcase the shortcomings of using LLMs as end-to-end customizers.
@@ 2.1 Task Formulation
Error Type
Goal $G$
Uncustomized Procedure $P$
Customization Hint $H$
Customized Procedure $P^{\prime}$
Wrong Order
Bake Bread
1. Dissolve yeast in water.
@@ 4. Put dough into greased pan
Customize Shoes
1. doodle on shoes.
@@
I am a ballet dancer and would like to improve the comfort of my shoes.
1. Identify areas of discomfort.
@@ 3.2 How Do The Agents Interact?
Figure 5: Modify agent and Verify agent in Parallel setting. Here, Modify agent suggests edits to customize the procedure $P$, while Verify agent suggests adding often-missing details so that $P$ can be followed. Finally, Resolve agent takes both sets of edits into account and produces a final bag of edits to be applied to $P$.
Parallel - In this setting, shown in Figure 5, both Modify agent and Verify agent propose a bag of edits for their respective aspect, $E_{c}$ and $E_{e}$, on the uncustomized procedure $P$. Since changes for customization and execution are different, conflicts arise between those bags of edits. It is non-trivial to understand how to select the correct edit, since both conflicting edits are important to generate $P^{\prime}$ but serve different objectives. To address this, we use Resolve agent , an LLM that is prompted to take as input two bags of edits, $E_{c}$ and $E_{e}$, and produce a merged set of edits $E$. Finally, applying $E$ on the uncustomized procedure $P$ results in $P^{\prime}$. Functionally, Resolve agent is intended to resolve conflicts, merge possible edits and remove redundant edits. It is also required to remove any edits which cannot be applied to the procedure deterministically. This is inspired by the meta-reasoner in Yoran et al. (2023).
@@ 4.2 Evaluation
We conduct a human evaluation with a standardized interface to compare different models. To this end, we pose questions about customizability, whether a procedure can be performed as is to accomplish the given goal (Executable), and executability, whether it satisfies all the requirements in a presented customization hint (Customized). This serves as the human evaluation interface for our task. The study is illustrated in Figure 7. The task instructions can be found in Appendix B.
@@ 4.3 Results
Resolving conflicts is difficult. We find that the Parallel setting is the worst, even though its performance on Customized and Executable are similar to other methods. However, the intersection of both (FullyCorrect) is significantly lower than others. We hypothesize that this problems arises because Resolve agent is unable to correctly merge bags of edits from Modify agent and Verify agent . This agent can be improved by providing a more complete outlook of how to resolve conflicts and merge relevant edits.
@@ 7 Conclusion
We acknowledge that there can be multiple ways to interpret a customization hint and only the user providing that hint can truly know their needs from the procedure. This also means that only that person is the right person to evaluate a customized procedure. Even so, due to the open-domain nature of this task, it is difficult for any one person to evaluate the various aspects of generated procedures. This can only be done by domain experts. To alleviate these problems, we use 3 annotators (master turkers) to judge model generated plans and consider their majority judgment, but recognize that this might not be the perfect solution.

### v3

@@ Abstract
Yash Kumar Lal13thanks: Work done as an intern at AI2 Aristo, Li Zhang23, Faeze Brahman3, Bodhisattwa Prasad Majumder3, Peter Clark3, Niket Tandon3 1 Stony Brook University, 2 University of Pennsylvania 3 Allen Institute for Artificial Intelligence 1ylal@cs.stonybrook.edu, 2zharry@upenn.edu, 3{faezeb,bodhisattwam,peterc,nikett}@allenai.org
@@ 2 Task Setup
In this section, we define a problem formulation and evaluation scheme for procedure customization and showcase the shortcomings of using LLMs as end-to-end customizers. LLMs are capable of generating fluent texts, including procedures Sakaguchi et al. (2021); Lyu et al. (2021). However, because LLMs generate texts in an autoregressive manner based on the previous context, they cannot edit those texts like humans would. This means that they need to rely on re-generation, which leads to unsatisfactory performance on our proposed task.
@@ 2.1 Task Formulation
Error Type Goal $G$ Uncustomized Procedure $P$ Customization Hint $H$ Customized Procedure $P^{\prime}$
Wrong Order Bake Bread 1. Dissolve yeast in water.
@@ 4. Put dough into greased pan
Customize Shoes 1. doodle on shoes.
@@
I am a ballet dancer and would like to improve the comfort of my shoes. 1. Identify areas of discomfort.
@@ 3.2 How Do The Agents Interact?
Figure 5: Modify agent and Verify agent in Parallel setting. Here, Modify agent suggests edits to customize the procedure $P$, while Verify agent suggests adding often-missing details so that $P$ can be followed. Finally, the edit resolution module takes both sets of edits into account and produces a final bag of edits to be applied to $P$.
Parallel - In this setting, shown in Figure 5, both Modify agent and Verify agent propose a bag of edits for their respective aspect, $E_{c}$ and $E_{e}$, on the uncustomized procedure $P$. Since changes for customization and execution are different, conflicts arise between those bags of edits. It is non-trivial to understand how to select the correct edit, since both conflicting edits are important to generate $P^{\prime}$ but serve different objectives. To address this, we use an edit resolution module, an LLM that is prompted to take as input two bags of edits, $E_{c}$ and $E_{e}$, and produce a merged set of edits $E$. Finally, applying $E$ on the uncustomized procedure $P$ results in $P^{\prime}$. Functionally, this module is intended to resolve conflicts, merge possible edits and remove redundant edits. It is also required to remove any edits which cannot be applied to the procedure deterministically. This is inspired by the meta-reasoner in Yoran et al. (2023).
@@ 4.2 Evaluation
We conduct a human evaluation with a standardized interface to compare different models. To this end, we pose questions about customizability, whether a procedure can be performed as is to accomplish the given goal (Executable), and executability, whether it satisfies all the requirements in a presented customization hint (Customized). This serves as the human evaluation interface for our task. The study is illustrated in Figure 7. The task instructions and annotator details can be found in Appendix B.
@@ 4.3 Results
Resolving conflicts is difficult. We find that the Parallel setting is the worst, even though its performance on Customized and Executable are similar to other methods. However, the intersection of both (FullyCorrect) is significantly lower than others. We hypothesize that this problems arises because the edit resolution module is unable to correctly merge bags of edits from Modify agent and Verify agent . This agent can be improved by providing a more complete outlook of how to resolve conflicts and merge relevant edits.
@@ 7 Conclusion
We acknowledge that there can be multiple ways to interpret a customization hint and only the user providing that hint can truly know their needs from the procedure. This also means that only that person is the right person to evaluate a customized procedure. Even so, due to the open-domain nature of this task, it is difficult for any one person to evaluate the various aspects of generated procedures. This can only be done by domain experts. To alleviate these problems, we use 3 annotators (master turkers) to judge model generated plans and consider their majority judgment Lal et al. (2022), but recognize that this might not be the perfect solution.
Due to the nature of LLMs, they are unable to encapsulate personal preferences V Ganesan et al. (2023); Dey et al. (2024). Since the procedure customization inherently involves user preferences, LLMs cannot be reliably used as automatic judges yet.